40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Message ID: 8923     Entry time: Thu Jul 25 13:54:35 2013
Author: manasa 
Type: Update 
Category: General 
Subject: PR3 clamped and Y arm is back flashing 

[Jenne, Annalisa, Manasa]

After yesterday's flipping of PR3, we lost our input pointing. Koji spent a few hours last night but couldn't restore the Y arm. I did my set of trials this morning which also didn't help.

So Jenne and I went ahead and requested Steve to get the ETMY door off.

We set the tiptilts TT1 and TT2 to the slider values from yesterday and started aligning the PR3 to hit the center of ITMY.
When we were hitting close to the center of ITMY, we decide to use the tip-tilts because the movement of PR3 was coarse at this point.
We used TT1 to get the beam to the center of ITMY and TT2 to get the beam at the center of ETMY. We did this iteratively until we were at the center of both the ITMY and ETMY.
We then went to fix IPANG.
The IPANG steering mirror on the BS table was steered to hit the center of the steering mirrors at the ETMY table. We aligned the beam to the IPANG QPD on the green endtable. The steering mirror on the BS table was then steered to misalign the beam in pitch by an inch at the last IPANG steering mirror. This should fix the IPANG clipping we have everytime we pump down.
We closed the chambers with light doors and saw IR flashing in the arm cavity. Koji is now trying to lock the cavity with IR.

ELOG V3.1.3-